OpenStudy (mathisfun13):

find the value of a1 for an infinite geometric series with S=-45 and r=-1/9.....i get -5 but its saying its wrong. idk how

jimthompson5910 (jim_thompson5910):

let a = first term S = a/(1-r) -45 = a/(1-(-1/9)) -45 = a/(1+1/9) ... ... ... a = ???

OpenStudy (mathisfun13):

.....4?

OpenStudy (mathisfun13):

oh wait

jimthompson5910 (jim_thompson5910):

let a = first term S = a/(1-r) -45 = a/(1-(-1/9)) -45 = a/(1+1/9) -45 = a/(9/9+1/9) -45 = a/(10/9) ... ... ... a = ???

OpenStudy (mathisfun13):

50

OpenStudy (mathisfun13):

-50

jimthompson5910 (jim_thompson5910):

let a = first term S = a/(1-r) -45 = a/(1-(-1/9)) -45 = a/(1+1/9) -45 = a/(9/9+1/9) -45 = a/(10/9) -45 = 9a/10 -45*10 = 9a ... ... ... a = ???

OpenStudy (mathisfun13):

i still get -50...

jimthompson5910 (jim_thompson5910):

good, wanted to make sure you weren't just guessing